18th July 2006 Observatory meeting room 11.30 - 13:35 Present: DWE, MJI, JRL, PSB, MR, STH, WJS Apologies: RGM, EGS Agenda ------ 1. Actions from last meeting 2. Comments on WFAU minutes 3. Data archives update - AAT, ING, WFCAM, UKIRT 4. Optical processing update 5. WFCAM news - reprocessing 05B - progress with 06A - non-linearity update - MSB ID checksum for tagging bad MSBs 6. VISTA news - VIRCAM AIT status and data - VDFS docs status - DRL status 7. AOB - hardware update - grants status - upcoming VDMT meeting - UKLight saga Minutes ------- 1. Actions from last meeting MJI has made no further progress investigating the merging of classification statistics. Most of the relevant information has been passed on, so it was decided to drop this action. JRL has started implemented the fix to the problem with the UKIRT archive (missing header information problem) - ongoing <<<< EGS has not updated the INT WFC web pages with the latest flat-field and fringe frames - ongoing <<<< STH has acquired the data from the 2MASS Touchstone fields. There are 7 fields that overlap our standard fields. It was thought that it would be a good idea to add this data to the VISTA Calibration Plan (if the data could be made public). STH was asked to update the Calibration Plan document in general. <<<< DWE are still finalising the content of Plone site. DWE said that many EGS of the technical pages had been written - ongoing <<<< STH JRL asked DWE to help with setting up a local ADASS Plone site MR (this was an action a few weeks back - but got lost) <<<< MJI arranged with RWA to fill in the 2004 gap for the INT in the EGS ING archive. This is now underway. JRL emailed MJI on the whereabouts of dome linearity sequences MR emailed Paul Hirst concerning incorrect offsets in the file headers EGS have started investigating the UDS claims regarding the pipeline MR components not being optimal. MR reported that he had received the data this morning (including the list of frames that made up the stack). 2. Comments on WFAU minutes JRL was a little puzzled by the comments regarding problems with the catalogue generating software since he had not received any emails saying that it wasn't working. He said that he would email WFAU to check out what the problem is. <<<< MJI said that there might be some subtle consequences caused by the changes between different versions of the software. MJI, MR and JRL will follow this up. <<<< CASU would like to congratulate WFAU for managing to get their site upgraded from 3 Amp round-pinned sockets :) It was pointed out that having a proactive specification for the air conditioning has meant that, even in heatwave conditions, we have enough capacity to keep our beer cold. 3. Data archives update AAT - no news ING - no news WFCAM - MR reported that tapes up to 5th July had been read onto disc and ingested into the raw data archive. ESO have transfered up to 10 June. The new data does not have any missing chips, but there are still a few files with zeroes in one of the chips. These tend to be at the start of the night. MJI said that a large proportion of the data from the first month was engineering data and was wondering if there was a way to stop this data being transfered to Cambridge. MJI will email JAC about this. <<<< UKIRT - no news 4. Optical processing update No news 5. WFCAM news Reprocessing 05B: MJI said that about 6 nights have been reprocessed using the latest sky subtraction algorithm. These mainly concerned UDS processing. The slow speed of feedback was regrettable. Progress with 06A: None of the data so far has been flagged as "ready to copy". The reason for this is that the Y and Z calibration needs a bit of tweaking due to complications arising from Galactic extinction. STH said that a Z comparison with respect to Sloan z' is being carried out. The bottom line is the the colour terms are within 0.5%. Non-linearity update: STH reported that the recently taken dome flats have revealed that WFCAM is linear to within 1%. The important thing is to track the variation in the background that is occurring while the dome flats are being taken. This is done by taking short exposures between each test exposure. There is no deviation above 1% until 40,000 counts, which is where saturation begins to take effect. This agrees with the previous dome flats. Also tests using standard star sequences show that WFCAM is linear up to saturation. It was noted that the linearity of 2MASS has only been tested to the 2% level. MSB ID checksum for tagging bad MSBs: JRL said that the JAC ID checksum method doesn't uniquely identify an MSB and is the reason why the information from the OMP database relating to bad MSBs has not been used. JRL and MJI will look into this further. <<<< It was noted that there were a few pipeline problems that needed fixing: Microstep sequence problems Header problems Deletion of XSky files (intermediate sky files) JRL said that he would get on these. <<<< 6. VISTA news VIRCAM AIT status and data: WJS said that the telescope was now in the dome and that the large external crane had now gone away. There will be an acceptance test at the end of August. The mirror timescale is in not so good a state. This will ship in late September with the secondary being OK, but there being problems with the primary. The camera was cold tested with an illuminated spot throughout June with PSB visiting RAL for one day. A current worry is that one of the chips has an area a little larger than a channel not working. This has been looked at under a microscope and no problem seen. If this is caused by a crack then the concern is that this will propagate and result in the entire chip not working. To alleviate this problem a science grade chip has been acquired. The camera has now been moved to the other end of the building close to the transport container. Some handling trials will be carried out next. It was unclear whether the camera will be going back into the clean room before it gets shipped. PSB said that another day visit to RAL to gather data would be nice. There is also the possibility of carrying out tests at the Paranal laboratory. The current schedule has the camera being shipped in late September. JRL said that he had started writing a report on the darks obtained from the AIT tests. The data has showed that the curtaining level is low (+/- 10 counts) and its removal is possible. Hot pixels do subtract out, except for chip 16. There is no crosstalk and there is persistence (more data needed). PSB said that he has had a request from Sandra Castro (?) to provide some benchmarks and that these were needed by the end of the month. VDFS docs status: PSB said that there was no change. MJI said that he would raise this issue at the VDMT meeting. <<<< WJS mentioned that the camera now has a new readout mode. This involves a read and a reset and is carried out on a per row basis and should be faster. (Destructive Read Mode?) It was pointed out that no modification can be carried out to the ETC until JPE provides CASU with the overhead algorithm. PSB said that he will be starting to digitize the graphs needed for the ETC help files (see minutes passim ad nauseam). DRL status: No news. 7. AOB Hardware update: MJI said that since we now have air conditioning in the storeroom, 2 of the more noisy machines have been moved into there. There is a need to sort out apm12 so that it can be included in the processing system. JRL will arrange with JMI to Debianize it. <<<< STH asked about the location of data. MR pointed him at the DQC database. Grants status: our funding is complex and is provided by several grants the CASU grant fully covers MJI, RWA, MR, STH and EGS until end March 2008 and partially covers the same folk until end March 2010 the CASU grant also includes our research activities until end March 2010 the e-science II grant which includes the VDFS development covers 1/2 PSB, DWE and JRL fully and 1/2 of STH and lasts until end September 2007 there is also the existing AstroGrid grant that covers 1/2 of EGS until end December 2006 the next grant due to start will be for Gaia and the majority of this starts in October 2007. With this, and some juggling of tasks, our current funding situation means that we can fund all the group until 2011. We also have some flexibility about how involved various group members want to be with Gaia. In addition there are also the new AstroGrid and VPHAS operations support grants that have just been submitted. This means that the group will expand by a few whatever else happens. The next big question is do we want to operate the VISTA pipeline? If we do this we will add another ~5 people to the group. We need to think soon if this is a desirable option and furthermore who is willing to manage this. Upcoming VDMT meeting: There is a VDMT meeting in Edinburgh on 24th July. STH pointed out that some people had been very naughty and not handed in their reports and that he wanted them in by tomorrow. This received the <<<< usual response. Additionally, STH wanted short reports on the AIT and the DRL from PSB and JRL respectively. <<<< MJI said that the UK VDFS Review will be held on 2 days in the week 23-27 October. We are to keep these dates free. The panel will be chaired by Luc Simart (sp?) and consist of Gavin Dalton, Tim Naylor, Matt Jarvis and Bob Nichol. It was pointed out that the Plone site will need to be clear as to the location of all our documentation. We will also need to know what documents we need to present at this review. UKLight saga: PSB asked whether we should go ahead with buying the hardware needed at our end. We all shouted YES. PSB was asked to go ahead with this. <<<< STH drew our attention to an email from SJW regarding "depreciations". He said that he would pass this on. <<<< STH said that there would be an UKIDSS meeting on 25th July in Edinburgh at which he and MJI would be giving presentations. STH also drew our attention to various papers that were in the process of being published/refereed. MR announced that he was getting married to Francesca on the 26th August. Many congratulations from CASU! Continuing Actions ------------------ JRL finish the fix to problem with UKIRT archive (missing header information problem) EGS update INT WFC web pages with latest flat-field and fringe frames DWE finalize content of Plone site by end of month (which one ?) EGS STH MR DWE help JRL in setting up an ADASS 2007 web site EGS New Actions ----------- STH update the Calibration Plan document (VIS-SPE-IOA-20000-0002 - check with PSB) JRL contact WFAU to clarify what is not working with the catalogue generating software. MJI compare old and new versions of software MR JRL MJI email JAC to see if it would be possible to exclude engineering data from the data sent to Cambridge JRL investigate MSB ID checksum for tagging bad MSBs MJI JRL fix pipeline problems (microstep, header, XSky) MJI raise issue of VDFS document status at VDMT meeting JRL arrange Debianizing apm12 with JMI (ALL ponder issue of CASU doing the VISTA reductions) ALL send in quarterly reports to STH (if not already done) PSB send in AIT and DRL report to STH JRL PSB set in motion acquiring UKLight hardware STH circulate SJW email regarding "depreciations"